A fluid motion is given by V =(y +z) i + (z + x) j +(x+ y) k. Show that themotion is irrotational and hence find the velocity potential also show that motion inpossible for an incompressible fluid.
1
Expert's answer
2020-10-29T17:47:57-0400

For steady incompressible flow, the co ntinuity equation is

["\partial u / \partial x + \partial v/\partial y+\partial w/\partial z" ]=0

u=y+z , v=z+x, w=x+y

"\partial(y+z)/\partial x+\partial(z+x)/\partial y+\partial(x+y)/\partial z=0\\
[0+0+0]=0\\
Therefore,the given flow field is a possible case of steady incompressible fluid flow\\" --show that the motion is irrotational

V=(y+z)i + (z+x)j +(x+y)k

V is irrotational if curl V=0

"I"


curl V="\nabla×V"

"\begin{vmatrix}
 i &&& j&&& k\\hi
 \partial/ \partial x &&& \partial/\partial y&&& \partial/ \partial z \\
(y+z)&&&(z+x)&&&(x+y)
\end{vmatrix}\\
i[\partial/\partial y(x+y)-\partial/\partial z (z+x)]-j[\partial/\partial x(x+y)-\partial/\partial z(y+z)]+k[\partial/\partial x(z+x)-\partial/ \partial y(y+z)]\\
=0\\
Curl V=\nabla×V=0\\
V \> is \>irrotational\\
3) Velocity \> potential=\phi =? \\
V=\nabla\phi------(1)\\
d\phi={(\partial\phi/\partial x)dx + (\partial\phi/\partial y)dy + (\partial\phi/\partial z)dz}.(idx +jdy+kdz)\\
d\phi=(i\partial /\partial x)+(j\partial/\partial y)+(k\partial/\partial z)\phi.dr \\
d\phi = \nabla \phi.dr------(2) \\
dr=idx+jdy+kdz \\
V=\nabla\phi \> put \> in \> eqn(2) \\
d\phi = V.dr \> equation (3)\\ 
V=(y+z)i+(z+x)j+(x+y)k \\
put\>in \> eqn(3) \\
d\phi={(y+z)i +(z+x)j + (x+y)k}.dr\\
d\phi={(y+z)i +(z+x)j + (x+y)k}.(idx+jdy+dz) \\
d\phi=(y+z)dx + (z+x)dy + (x+y)dz \>(4)\\
d\phi = ydx+zdx+zdy+xdy+xdz+ydz\\
d\phi=(xdy+ydx)+(xdz+zdx)+(ydz+zdy)\\
Product \> of \> two \> function\\
d\phi=dxy+dxz+dyz\\
On \> integration \\
\phi=xy+xz+yz+ C \\
Where\>C\> is \> integration \> constant\\
\phi \> is\> the\>velocity\>potential."